The length of a rectangle is (x+2) inches. The width of the rectangle is ( x + 1) inches. Write an equation that can be used to determine the area (A) of the rectangle in terms of x. What is the value of x is the area of the rectangle is 56 square inches?
Answer by edjones(8007)

A=(x+1)(x+2)=x^2+3x+2
x^2+3x+2=56
x^2+3x-54=0
(x-6)(x+9)=0
x=6"
.
Ed